OpenStudy (anonymous):

explain the difference between 1/(sinx) and ((sin^-1)x)

OpenStudy (anonymous):

\[\frac{1}{sin(x)}=csc(x)\] \[sin^{-1}(x)=arcsin(x)\]

OpenStudy (anonymous):

\(1/\sin x\) is the inverse value of \(\sin x\). \(\sin^{-1} x\) is the inverse function of sine (of \(x\)).

OpenStudy (anonymous):

Sin(pi/6)=1/2 ArcSin(1/2)=pi/6

OpenStudy (anonymous):

Personally, I would always say arcsin if that's what I meant, the other can cause confusion.

OpenStudy (anonymous):

Writing \(\sin^{-1}\) is actually very systematic since it follows the general composite function notation. \(\sin^{-1}\) is an equivalent marking to that of \(f^{-1}\), denoting the inverse function. "Going back one step." \(\sin\) just happens to be a familiar function with a name we all recognise. The notation \(\sin^{2}x\) again is misleading since it doesn't mean \(\sin(\sin(x)))\) but \((\sin x)^2\).

OpenStudy (anonymous):

I look at it precisely in the reverse way, though (so do programmers).